Make New Entry,
Make Followup Entry
User name paschke
Log entry time 04:36:38 on August24,2009
Entry number 285631
keyword=improved carbon spectrum in left arm after S0 swap
Carbon spectra on the left arm suggest that the problem was in the efficency of S0.
First attachment: Run 24511, Sieve in -3% on carbon. 20 uA. S0 Counter 1.
Second attachment: Run 24486, Sieve in, -3% on carbon. 20 uA, S0 Counter 2 (now in right arm).